OpenStudy (anonymous):

Please explain step by step... Piare bought a self-assembly desk and had to use a wrench to tighten nuts he tried that one and a half inch wrench which was too small and then nine sixteent inch wrench which was too large. Which metric wrench should he try.. A. 10mm B. 12mm C. 14mm D. 15mm

OpenStudy (triciaal):

2,5 cm = 1 inch 10mm = 1 cm (25mm = 1 inch) 9/16 x 2.5 = cm convert to mm this is too large 1 and 1/2 inch = 3/2 x 2.5cm convert to mm too small the correct size would be between the one that is too small and the one that is too large. I think you wrote something wrong you have 9/16 is more than one and a half please check.